OpenStudy (anonymous):

factor x2 + 4/3 - 4/9

OpenStudy (anonymous):

do you mean x^2+4/3x-4/9?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you may try to solve for x using quadratic formula on x^2+4/3x-4/9=0.

OpenStudy (theviper):

\[\Huge{\color{red}{x^2+{8\over9}}}\]

OpenStudy (anonymous):

then, after that, (x - value you got)(x+value you got).

OpenStudy (anonymous):

that is, if the original problem was x^2+4/3x-4/9.

OpenStudy (anonymous):

@TheViper how you get it ?

OpenStudy (theviper):

Solve it

OpenStudy (theviper):

Subtract 4/3 from 4/9,i.e., \[\Huge{\color{green}{\frac{4}{3}-\frac{4}{9}}}\]

OpenStudy (anonymous):

janirish said that the original problem was x^2+4/3x-4/9.

OpenStudy (anonymous):

the factor of this will have a radical, and the best method to use is quadratic formula.

OpenStudy (theviper):

ok ok then @violy is correct:)

OpenStudy (anonymous):

ok ! now what is the correct answer ?

OpenStudy (anonymous):

try to use the quadratic formula